OpenStudy (anonymous):

(4x)3 4x3 7x3 12x3 64x3

OpenStudy (bibby):

\(\large (ab)^c=a^b*a^c\)

OpenStudy (anonymous):

i dont understand @bibby

OpenStudy (bibby):

for example \((6b^2)^7=6^7*({b^2})^7=6^7*b^{14}\)

OpenStudy (anonymous):

so would it be (4x)3= 4^3 x (x^1) ? @bibby

OpenStudy (bibby):

x^3*

OpenStudy (bibby):

the point is the power gets distributed to each term being multiplied

OpenStudy (bibby):